# Scaling Latent Reasoning via Looped Language Models

### v3

@@ 3 Learning Adaptive Latent Reasoning with LoopLM
Figure 3: Overview of Looped Language Model (LoopLM) architecture. Left (Training): During training, the model applies a stack of NN layers repeatedly for $T_{max}$ recurrent steps. At each recurrent step $\ell$, an exit gate predicts the probability $p_{\ell}$ of exiting, and a language modeling head $\mathcal{L}_{\ell}$ computes the lanugage modeling loss. Right (Inference): At inference time, the model can exit early based on the accumulated exit probability.
@@ 3.1 LoopLM Architecture
where $p^{(t)}_{\theta}(\cdot\mid x_{1:\ell})=\mathrm{softmax}\!\big(\mathrm{lmhead}(h^{(t)}_{\ell})\big)$, $x_{1:\ell}$ denotes the length-$\ell$ prefix of the input (tokens 11 through ellell), and $h^{(t)}_{\ell}$ is the hidden state after $t$ loops at position $\ell$. Note that this is the individual loss for a single recurrent step. The total training objective, which combines all steps, is defined in the following sections.
@@ 4.2 Data
Nemotron-$C$ (Web Data) Stage 1 6386 4404
MAP-$C$ (Web Data) Stage 1 800 780
@@
This stage supports the warmup and stable phases of training. The corpus is primarily composed of Web CommonCrawl ($C$) data. Because we sought to train the model on >2T tokens, many popular open corpora are too small (e.g., Fineweb-Edu at 1.3T tokens [38], DCLM at 2.6T tokens [39]). We therefore use Nemotron-$C$ [40] (6.3T tokens) as the main dataset for the stable-phase. To provide the model with basic Chinese proficiency, we include Ultra-FineWeb-zh [41] and MAP-$C$ [42]. However, without Chinese vocabulary in the tokenizer, characters would be fragmented into multiple byte-level sub-tokens, so we removed Chinese from Stage 2 onwards. To enhance coding and mathematical abilities, we incorporate OpenCoder [43] and MegaMath [44]. See Table˜4 for dataset proportions in further detail.
@@
The CT annealing stage incorporates higher-quality data to enhance the model under the annealing learning rate. Token sequence length is extended to 16K tokens, exceeding the length of most samples to minimize truncation. We construct the corpus from the high-quality subset of Nemotron-$C$ and augment with HQ MegaMath, Nemotron-$C$-Math-v1 [45, 46], OpenCoder-Annealing [43], Nemotron-pre-training-Code-v1 [46], and Nemotron-pre-training-SFT-v1 [46]. Data composition is provided inTable˜5.
@@ 6.1 LoopLMs does not increase knowledge capacity
Settings. Following the Capo task setting in Physics of language models [67, 68], we construct synthetic biographies to test how much information the model memorizes. Specifically, we generate several synthetic biographic datasets $\operatorname{bioS}(N)$ with different number of people NN, and train a series of language models to memorize the information contained in the dataset. Each biography contains the individual’s name and five attributes $a_{1},a_{2},...,a_{5}$ of the person: gender, birth date, university, major, and employer. The names $n$ and the attributes $a_{i}$ are randomly selected from a pre-defined set $\mathcal{N}$ and $\mathcal{A}_{i}$ and combined together as a biography using a random template. Based on the random generation process, we have an information-theoretic lower bound for the model in the minimum bits required to encode all the names and attributes. To check whether the models memorize the biographic information accurately, we look at the probability of predicting the ground-truth attributes with the trained models given the biography context. Calculating the sum of cross-entropy loss on each attribute token positions, we can estimate how much information (estimated in bits) has already been memorized in the trained language model, which is our knowledge capacity metric.
With this metric, we can compare the knowledge capacity between the original model (with only one recurrent step) and the looped model (with 4 recurrent steps) with the same parameter count to investigate whether looping increases knowledge capacity. Moreover, as larger models should encode more information than smaller models, we also aim to investigate whether looped models have a better scaling effect when the size of the model grows. We thereby trained GPT-2 style models of different parameter numbers ranging from 1M to 40M (with depth and hidden dimension varied) and measured the number of bits of knowledge learned by each model. We trained on $\operatorname{bioS}(N)$ with NN ranging from 20K to 500K individuals for 1000 exposures. More training details are provided in Section˜B.1.
@@
Figure 6: Left..
@@ 7.2 Faithfulness
In our model, the situation is very different. our $1.4\mathrm{B}\!\times\!4$ model uses 24 layers per recurrent step. We train linear probes on hidden states from layers 11 through $24i$ to predict the step-$i$ answer, for $i\!\in\!\{2,3,4\}$. Within a single recurrent step, the step-$i$ answer is well predicted by a probe on representation within layer $24i$, indicating strong intra-step alignment between state and decision, which is similar to the non-reasoning model Qwen-4B-Instruct, showing in left part of Figure 9. Crucially, probes on the preceding representation (layer 24(i1)24(i\!-\!1)) do not reliably predict the step-$i$ decision for $i\!\in\!\{2,3,4\}$, showing that the new recurrent pass performs additional computation that can revise a provisional choice.
To further examine the consistency between the results of different rounds. We also compute a step-by-step agreement matrix AA over 1,000 Quora Question Pairs, where $A[i,j]$ counts identical labels between step $i$ and step $j$ (diagonal =1000=\!1000 by construction). See the right side of Figure 9. Adjacent steps never reach full agreement; for example, A[2,4]=361A[2,4]\!=\!361 indicates only $36.1\%$ of step-2 answers match step-4. A[2,3]=551A[2,3]\!=\!551 indicates only $55.1\%$ of step-2 answers match step-3. We also notice that when $i\geq 4$, the overlap consistency between step-$i$ and step-$i+1$, $A[i,i+1]$, is close to 1000. We think this phenomenon comes from: (1) the model does not learn to reason recursively when $i>4$. The model is trained within 4 loops; (2) as the number of loops increases, the answer gradually converges to a fixed point.
@@ B.1 Capo: knowledge capacity
We synthesize fake biographies following the $\mathrm{bioS}(N)$ dataset in [67]. Specifically, we generate NN biographies of a random generated person together with their date of birth, city of birth, university, major, and employer. In our work, we online sample the individual attributes and generate the biographies in natural language using a random selected fixed template. An illustrative example is:
@@
After pre-training on the bioS(NN) dataset, we assess a model’s knowledge capacity, defined as the number of bits of information it can reliably store. To make this measure comparable across models of different sizes, the raw bit count is normalized by the number of model parameters, yielding a “bits per parameter” metric. The derivation and motivation of the metric in discussed in [67]. For readers, we refer the detailed setting to Section 2.1 of [67].
@@
Given a model $F$ with $P$ parameters trained over the bioS(NN) dataset $Z$, suppose it gives $p_{1}=\text{loss}_{\text{name}}(Z)$ and $p_{2}=\text{loss}_{\text{value}}(Z)$, which are the sum of cross entropy loss on the name tokens and attribute tokens, respectively. The capacity ratio and the maximum achievable capacity ratio are defined as
@@ B.3 Multi-hop question answering on synthetic relations
The dataset contains $|\mathcal{E}|$ entities—each with a unique name—and NN relation types. We created 500 distinct single-token person names (e.g., Jennifer) and 20 single-token relation names (e.g., instructor) to serve as namespaces for entities and relations. We reused the name list in [69]. The complete list of relation names and a partial list of entity names appear in Tables 5 and 6 in [69]. The multi-hop questions are generated through a $K=5$ hierarchical layers, where each layer has 100 individuals. Each entity is connected to $|\mathcal{R}|$ randomly chosen person in the next layer. This structure naturally generates $|\mathcal{E}|/5\times|\mathcal{R}|^{k}$ $k$-hop questions. In our setting, since we only consider 3-hop questions, the number should be $8\times 10^{5}$.
@@ B.3.1 Additional experimental results
Figure 11: Left & Right. We further train with 100000100000 and 140000140000 unique QA pairs for 20000 steps with context length 1024 and batch size 2048. Similar to the main text, models with more loops learn faster and achieve better performance comparing with models without loops.
Figure 12: Left & Right. We further train with 100000100000 and 120000120000 unique QA pairs for 20000 steps with context length 1024 and batch size 2048. We train the baseline with 24 layers, which is equivalent flops with the loop 4 transformers. Similar to the main text, models with more loops learn faster and achieve better performance comparing with models without loops, even with iso-flop transformers. The loop 2 average performance is weaker than the iso-flop version transformer since it has less equivalent depth when $N=10^{5}$, but it surpasses the baseline with more data provided.
@@ B.5 Theory: latent thought with LoopLM
Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$ is the set of vertices and $E=\{e_{1},e_{2},\ldots,e_{m}\}$ is the set of edges. Let $G=(V,E)$ be a directed hidden knowledge graph, and $G_{\text{ctx}}=(V,E_{\text{ctx}})$ be an input additional knowledge graph. Given a source node $s$ a target node $t$, the task is to output 11 when there exists a path from $s$ to $t$ on the combined graph $G+G_{\text{ctx}}:=(V,E+E_{\text{ctx}})$, and output 0 when $s$ cannot reach $t$ on the combined graph.
@@ D.2 RQ2: How do recurrent step impact the total loss and step-wise loss in the context of LoopLM?
Here, we have following variables: model size NN, training data size $D$, maximum recurrent step $T_{m}$, recurrent step $T$, total loss $L_{t}$, and step-wise loss $L_{s}$. Following Chinchilla [96], we first attempt to fit the relationship between $L_{t}$ and $N,D,T_{m}$ in the form of a power law:
@@
In addition to its predictability, we further explore the generalizability of the Total Loss Scaling Law. Predictability refers to the ability of the Scaling Law to fit all data points into a unified curve when all data points are available. Generalizability, on the other hand, indicates whether the Scaling Law can predict unseen data points when fitting is done with a subset of data points. For example, generalizability tests whether the performance of a 14B model can be predicted using the known performances of 1B and 7B models [97]. To verify its generalizability across model size NN, training data size $D$, and maximum recurrent step $T_{m}$, we have conducted related experiments, details can be found in Appendix E.1.
During the LoopLM training process, we compute the cross-entropy loss at each recurrent step, which we refer to as step-wise loss $L_{s}$. We aim to explore the relationship between step-wise loss $L_{s}$ and the current recurrent step $T$, model size NN, and training data size $D$. Similarly, we can fit the scaling law between $L_{s}$ and $N,D,T$, with the formula as follows:
@@ D.3 RQ3: What is the inherent connection between total loss and step-wise loss?
We further demonstrate this through practical experiments. We take the situation where the max recurrent step is equal to 4. First, we perform a standard fitting of the Step-wise Loss Scaling Law to obtain the fitting parameters $E_{s},A_{s},B_{s},C_{s},$ and so on. Next, we observe and record the distribution of $q_{\phi}$ for each recurrent step $t$ when the maximum recurrent step $T_{m}=4$, as shown in Figure 19. For convenience, we take the average value of $q_{\phi}$ at different recurrent steps and treat it as a normal discrete distribution, resulting in the distribution {0.0004, 0.0855, 0.3793, 0.5348}. We then substitute this distribution and the $T$ values into the training objective, ignoring the entropy regularization term (after the training stabilizes, it becomes relatively low, for simplicity, we will just ignore it). This leads to a fitting formula, and upon substituting the actual fitting data points NN and $D$, the computed $R^{2}$ value is 0.961, with the fitting results illustrated in Figure 20. We can see that the fitting accuracy is high, and the predicted curve closely matches the actual curve, indicating that, under a relatively rough estimate, step-wise loss can be transformed into total loss, thus indirectly suggesting an intrinsic connection between the two.

### v4

@@ 3 Learning Adaptive Latent Reasoning with LoopLM
Figure 3: Overview of Looped Language Model (LoopLM) architecture. Left (Training): During training, the model applies a stack of $N$ layers repeatedly for $T_{max}$ recurrent steps. At each recurrent step $\ell$, an exit gate predicts the probability $p_{\ell}$ of exiting, and a language modeling head $\mathcal{L}_{\ell}$ computes the lanugage modeling loss. Right (Inference): At inference time, the model can exit early based on the accumulated exit probability.
@@ 3.1 LoopLM Architecture
where $p^{(t)}_{\theta}(\cdot\mid x_{1:\ell})=\mathrm{softmax}\!\big(\mathrm{lmhead}(h^{(t)}_{\ell})\big)$, $x_{1:\ell}$ denotes the length-$\ell$ prefix of the input (tokens $1$ through ellell), and $h^{(t)}_{\ell}$ is the hidden state after $t$ loops at position $\ell$. Note that this is the individual loss for a single recurrent step. The total training objective, which combines all steps, is defined in the following sections.
@@ 4.2 Data
Nemotron-CC (Web Data) Stage 1 6386 4404
MAP-CC (Web Data) Stage 1 800 780
@@
This stage supports the warmup and stable phases of training. The corpus is primarily composed of Web CommonCrawl (CC) data. Because we sought to train the model on >2T tokens, many popular open corpora are too small (e.g., Fineweb-Edu at 1.3T tokens [38], DCLM at 2.6T tokens [39]). We therefore use Nemotron-CC [40] (6.3T tokens) as the main dataset for the stable-phase. To provide the model with basic Chinese proficiency, we include Ultra-FineWeb-zh [41] and MAP-CC [42]. However, without Chinese vocabulary in the tokenizer, characters would be fragmented into multiple byte-level sub-tokens, so we removed Chinese from Stage 2 onwards. To enhance coding and mathematical abilities, we incorporate OpenCoder [43] and MegaMath [44]. See Table˜4 for dataset proportions in further detail.
@@
The CT annealing stage incorporates higher-quality data to enhance the model under the annealing learning rate. Token sequence length is extended to 16K tokens, exceeding the length of most samples to minimize truncation. We construct the corpus from the high-quality subset of Nemotron-CC and augment with HQ MegaMath, Nemotron-CC-Math-v1 [45, 46], OpenCoder-Annealing [43], Nemotron-pre-training-Code-v1 [46], and Nemotron-pre-training-SFT-v1 [46]. Data composition is provided inTable˜5.
@@ 6.1 LoopLMs does not increase knowledge capacity
Settings. Following the Capo task setting in Physics of language models [67, 68], we construct synthetic biographies to test how much information the model memorizes. Specifically, we generate several synthetic biographic datasets $\operatorname{bioS}(N)$ with different number of people $N$, and train a series of language models to memorize the information contained in the dataset. Each biography contains the individual’s name and five attributes $a_{1},a_{2},...,a_{5}$ of the person: gender, birth date, university, major, and employer. The names $n$ and the attributes $a_{i}$ are randomly selected from a pre-defined set $\mathcal{N}$ and $\mathcal{A}_{i}$ and combined together as a biography using a random template. Based on the random generation process, we have an information-theoretic lower bound for the model in the minimum bits required to encode all the names and attributes. To check whether the models memorize the biographic information accurately, we look at the probability of predicting the ground-truth attributes with the trained models given the biography context. Calculating the sum of cross-entropy loss on each attribute token positions, we can estimate how much information (estimated in bits) has already been memorized in the trained language model, which is our knowledge capacity metric.
With this metric, we can compare the knowledge capacity between the original model (with only one recurrent step) and the looped model (with 4 recurrent steps) with the same parameter count to investigate whether looping increases knowledge capacity. Moreover, as larger models should encode more information than smaller models, we also aim to investigate whether looped models have a better scaling effect when the size of the model grows. We thereby trained GPT-2 style models of different parameter numbers ranging from 1M to 40M (with depth and hidden dimension varied) and measured the number of bits of knowledge learned by each model. We trained on $\operatorname{bioS}(N)$ with $N$ ranging from 20K to 500K individuals for 1000 exposures. More training details are provided in Section˜B.1.
@@
Figure 6: Left.We trained both LoopLM and a standard trasnformer baseline with the same parameters on Capo task to compare the knowledge capacity gain by looping more times. With the same parameter count, the looped model and its non-looped baseline has almost the same knowledge capacity measured in bits of knowledge on Capo task. Right. Accuracy of looped/non-looped models on Mano task. Looped models are better than the iso-param $(\{2,3,6\}\otimes 1)$ models. They also achieve better or comparable performance comparing to the iso-flop baseline $(12\otimes 1)$ model.
@@ 7.2 Faithfulness
In our model, the situation is very different. our $1.4\mathrm{B}\!\times\!4$ model uses 24 layers per recurrent step. We train linear probes on hidden states from layers $1$ through $24i$ to predict the step-$i$ answer, for $i\!\in\!\{2,3,4\}$. Within a single recurrent step, the step-$i$ answer is well predicted by a probe on representation within layer $24i$, indicating strong intra-step alignment between state and decision, which is similar to the non-reasoning model Qwen-4B-Instruct, showing in left part of Figure 9. Crucially, probes on the preceding representation (layer 24(i1)24(i\!-\!1)) do not reliably predict the step-$i$ decision for $i\!\in\!\{2,3,4\}$, showing that the new recurrent pass performs additional computation that can revise a provisional choice.
To further examine the consistency between the results of different rounds. We also compute a step-by-step agreement matrix $A$ over 1,000 Quora Question Pairs, where $A[i,j]$ counts identical labels between step $i$ and step $j$ (diagonal =1000=\!1000 by construction). See the right side of Figure 9. Adjacent steps never reach full agreement; for example, A[2,4]=361A[2,4]\!=\!361 indicates only $36.1\%$ of step-2 answers match step-4. A[2,3]=551A[2,3]\!=\!551 indicates only $55.1\%$ of step-2 answers match step-3. We also notice that when $i\geq 4$, the overlap consistency between step-$i$ and step-$i+1$, $A[i,i+1]$, is close to 1000. We think this phenomenon comes from: (1) the model does not learn to reason recursively when $i>4$. The model is trained within 4 loops; (2) as the number of loops increases, the answer gradually converges to a fixed point.
@@ B.1 Capo: knowledge capacity
We synthesize fake biographies following the $\mathrm{bioS}(N)$ dataset in [67]. Specifically, we generate $N$ biographies of a random generated person together with their date of birth, city of birth, university, major, and employer. In our work, we online sample the individual attributes and generate the biographies in natural language using a random selected fixed template. An illustrative example is:
@@
After pre-training on the bioS($N$) dataset, we assess a model’s knowledge capacity, defined as the number of bits of information it can reliably store. To make this measure comparable across models of different sizes, the raw bit count is normalized by the number of model parameters, yielding a “bits per parameter” metric. The derivation and motivation of the metric in discussed in [67]. For readers, we refer the detailed setting to Section 2.1 of [67].
@@
Given a model $F$ with $P$ parameters trained over the bioS($N$) dataset $Z$, suppose it gives $p_{1}=\text{loss}_{\text{name}}(Z)$ and $p_{2}=\text{loss}_{\text{value}}(Z)$, which are the sum of cross entropy loss on the name tokens and attribute tokens, respectively. The capacity ratio and the maximum achievable capacity ratio are defined as
@@ B.3 Multi-hop question answering on synthetic relations
The dataset contains $|\mathcal{E}|$ entities—each with a unique name—and $N$ relation types. We created 500 distinct single-token person names (e.g., Jennifer) and 20 single-token relation names (e.g., instructor) to serve as namespaces for entities and relations. We reused the name list in [69]. The complete list of relation names and a partial list of entity names appear in Tables 5 and 6 in [69]. The multi-hop questions are generated through a $K=5$ hierarchical layers, where each layer has 100 individuals. Each entity is connected to $|\mathcal{R}|$ randomly chosen person in the next layer. This structure naturally generates $|\mathcal{E}|/5\times|\mathcal{R}|^{k}$ $k$-hop questions. In our setting, since we only consider 3-hop questions, the number should be $8\times 10^{5}$.
@@ B.3.1 Additional experimental results
Figure 11: Left & Right. We further train with $100000$ and 140000140000 unique QA pairs for 20000 steps with context length 1024 and batch size 2048. Similar to the main text, models with more loops learn faster and achieve better performance comparing with models without loops.
Figure 12: Left & Right. We further train with $100000$ and 120000120000 unique QA pairs for 20000 steps with context length 1024 and batch size 2048. We train the baseline with 24 layers, which is equivalent flops with the loop 4 transformers. Similar to the main text, models with more loops learn faster and achieve better performance comparing with models without loops, even with iso-flop transformers. The loop 2 average performance is weaker than the iso-flop version transformer since it has less equivalent depth when $N=10^{5}$, but it surpasses the baseline with more data provided.
@@ B.5 Theory: latent thought with LoopLM
Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$ is the set of vertices and $E=\{e_{1},e_{2},\ldots,e_{m}\}$ is the set of edges. Let $G=(V,E)$ be a directed hidden knowledge graph, and $G_{\text{ctx}}=(V,E_{\text{ctx}})$ be an input additional knowledge graph. Given a source node $s$ a target node $t$, the task is to output $1$ when there exists a path from $s$ to $t$ on the combined graph $G+G_{\text{ctx}}:=(V,E+E_{\text{ctx}})$, and output 0 when $s$ cannot reach $t$ on the combined graph.
@@ D.2 RQ2: How do recurrent step impact the total loss and step-wise loss in the context of LoopLM?
Here, we have following variables: model size $N$, training data size $D$, maximum recurrent step $T_{m}$, recurrent step $T$, total loss $L_{t}$, and step-wise loss $L_{s}$. Following Chinchilla [96], we first attempt to fit the relationship between $L_{t}$ and $N,D,T_{m}$ in the form of a power law:
@@
In addition to its predictability, we further explore the generalizability of the Total Loss Scaling Law. Predictability refers to the ability of the Scaling Law to fit all data points into a unified curve when all data points are available. Generalizability, on the other hand, indicates whether the Scaling Law can predict unseen data points when fitting is done with a subset of data points. For example, generalizability tests whether the performance of a 14B model can be predicted using the known performances of 1B and 7B models [97]. To verify its generalizability across model size $N$, training data size $D$, and maximum recurrent step $T_{m}$, we have conducted related experiments, details can be found in Appendix E.1.
During the LoopLM training process, we compute the cross-entropy loss at each recurrent step, which we refer to as step-wise loss $L_{s}$. We aim to explore the relationship between step-wise loss $L_{s}$ and the current recurrent step $T$, model size $N$, and training data size $D$. Similarly, we can fit the scaling law between $L_{s}$ and $N,D,T$, with the formula as follows:
@@ D.3 RQ3: What is the inherent connection between total loss and step-wise loss?
We further demonstrate this through practical experiments. We take the situation where the max recurrent step is equal to 4. First, we perform a standard fitting of the Step-wise Loss Scaling Law to obtain the fitting parameters $E_{s},A_{s},B_{s},C_{s},$ and so on. Next, we observe and record the distribution of $q_{\phi}$ for each recurrent step $t$ when the maximum recurrent step $T_{m}=4$, as shown in Figure 19. For convenience, we take the average value of $q_{\phi}$ at different recurrent steps and treat it as a normal discrete distribution, resulting in the distribution {0.0004, 0.0855, 0.3793, 0.5348}. We then substitute this distribution and the $T$ values into the training objective, ignoring the entropy regularization term (after the training stabilizes, it becomes relatively low, for simplicity, we will just ignore it). This leads to a fitting formula, and upon substituting the actual fitting data points $N$ and $D$, the computed $R^{2}$ value is 0.961, with the fitting results illustrated in Figure 20. We can see that the fitting accuracy is high, and the predicted curve closely matches the actual curve, indicating that, under a relatively rough estimate, step-wise loss can be transformed into total loss, thus indirectly suggesting an intrinsic connection between the two.